OpenStudy (anonymous):

Calculate the discriminant. x2 – 2x + 4 = 0 A. 12 B. –12 C. 20 D. 0

OpenStudy (campbell_st):

the discriminant for a quadratic \[ax^2 + bx + c = 0\] is \[x = b^2 - 4 \times a \times c\] in the question a = 1, b = -2 and c = 4 so substitute and calculate the answer.... take care with squaring the negative

OpenStudy (accessdenied):

It might be more appropriate to use a different variable than x for the exact discriminant. Usually I see capital delta, but any old letter (d, D, etc.) would be fine: \(\color{blue}\Delta = b^2 - 4ac\) Since its motivation for quadratic equations stems from the quadratic formula: \( x = \dfrac{-b \pm \sqrt{\color{blue}{b^2 - 4ac}}}{2a} \qquad = \dfrac{-b \pm \sqrt{\color{blue}{\Delta}}}{2a}\) But regardless, calculating \( b^2 - 4ac \) from the equation is our goal. :)